# arXiv:2603.02657v1[cs.RO] 03 Mar 2026

### v1

@@ Abstract
Although legged robots demonstrate impressive mobility on rough terrain, using them safely in cluttered environments remains a challenge. A key issue is their inability to avoid stepping on low-lying objects, such as high-cost small devices or cables on flat ground. This limitation arises from a disconnection between high-level semantic understanding and low-level control, combined with errors in elevation maps during real-world operation. To address this, we introduce SemLoco, a Reinforcement Learning (RL) framework designed to avoid obstacles precisely in densely cluttered environments. SemLoco uses a two-stage RL approach that combines both soft and hard constraints. It performs pixel-wise foothold safety inference, which enables more accurate foot placement. Additionally, SemLoco integrates semantic map, allowing it to assign traversability costs instead of relying only on geometric data. SemLoco greatly reduces collisions and improves safety around sensitive objects, enabling reliable navigation in situations where traditional controllers would likely cause damage. Experimental results further show that SemLoco can be effectively applied to more complex, unstructured real-world environments.
@@ I Introduction
  • SemLoco Framework: We propose the SemLoco framework, which solves the exploration trap in the navigation of dense obstacles. By decoupling spatial exploration (via virtual obstacles) from strict kinematic tracking (via rigid physical interactions), our curriculum enables the policy to establish robust perception-motor mappings, thereby significantly enhancing foothold precision in highly cluttered spaces.
  • Semantic-Geometry Decoupling: We introduce a semantic-aware control pipeline that explicitly disentangles environmental semantics from geometric heights. This paradigm overcomes the inherent limitations of traditional elevation-only proxies while remaining highly robust against the inherent noise of physical depth sensors.
  • Sim-to-Real Deployment and Validation: We rigorously validate the proposed method across both simulations and real-world scenarios featuring fragile equipment. Extensive experiments demonstrate that SemLoco drastically reduces the step collision rate, proving its robust autonomy and adaptability in unstructured, real-world environments.
@@ II Related Works
Semantic Perception in Robot Navigation: Beyond purely geometry-based navigation, recent works leverage semantic information to enhance planning. Several modular approaches employ semantic segmentation models to obtain object masks and assign heuristic costs to construct dense semantic maps [3, 17] or semantic graphs [13]. [12] learns a semantic traversability estimator through an automated annotation pipeline on egocentric videos, simplifying the derivation of traversal cost maps. Furthermore, semantic-aware planners are built upon these representations to incorporate object-level risk into navigation decisions [2, 1]. More recently, [24] proposes an end-to-end path planning framework that encodes semantic information in the latent space, improving planning efficiency. [29] integrates semantic semantics into quadrupedal navigation, yet its impact is confined to the tuning of gait parameters, and velocity levels. These semantic-aware methods primarily operate at the high-level path planning stage and do not explicitly reason about foothold selection for legged robots. In cluttered and narrow environments, even if fragile or undesirable objects are recognized by the high-level planner, the absence of foothold-level constraints may still cause the robot to step on them, limiting the effectiveness of semantic awareness in contact-rich locomotion. To address the above issues, SemLoco  for the first time, combines semantic information with foothold planning for legged robots, using it to assign traversability costs to different types of obstacles, thereby establishing an effective connection between high-level semantic understanding and low-level motion control.
@@ III Methodology
Figure 2: Framework of SemLoco. Sub-modules have different styles based on their functions. Among them, the red trapezoid represents the neural network, the blue rectangle represents unprocessed raw data, and the green rectangle represents processed ready-to-use data. (a) Training in the simulator: In stage 1, we use virtual obstacles (Highly yellow spheres). Although the robot walks on flat ground, it receives a virtual perception map containing height and semantic information to simulate scenarios with real obstacles. In stage 2, we use rigid obstacles. The robot receives corresponding real perception information, and fine-tunes the policy to improve task performance. (b) Deployment in the real world: The semantic algorithm uses Odin1 to process exteroception. Combined with proprioception, this information is fed into the trained policy to generate low-level control commands.
@@ III-A1 Observation Space
where $\bm{C}_{t}^{\text{cmd}}\in\mathbb{R}^{1}$ are the desired velocity commands, including the base linear velocities $\bm{v}_{x}^{\text{cmd}}\in\mathbb{R}^{1}$ in the x-axes, $\bm{v}_{y}^{\text{cmd}}\in\mathbb{R}^{1}$ in the y-axes and the base angular velocity $\bm{\omega}_{z}^{\text{cmd}}\in\mathbb{R}^{1}$ in the yaw axis.
@@
The $\bm{O}_{t}^{\text{extero}}\in\mathbb{R}^{1440}$ are defined as information about exteroception. Environmental information includes the elevation map $\bm{H}_{t}^{\text{elev}}\in\mathbb{R}^{720}$ and the semantic map $\bm{H}_{t}^{\text{sem}}\in\mathbb{R}^{720}$, both of which are centered at the center of gravity of the robot. Both maps are defined over a region that spans 1.5 m in the x-direction and 1.2 m in the direction . With a spatial resolution of 0.05 m, this results in a 30 $\times$ 24 grid of sampled cells.
@@ III-A3 Network Architecture
The framework is illustrated as Fig. 2. Our policy consists of two Multilayer Perceptrons (MLPs) for the actor network and the critic network respectively, a base velocity estimator and a Convolutional Neural Network (CNN) encoder. The CNN encoder extracts height features from elevation map and semantic features from semantic map. The feature vector extracted by the CNN module is concatenated with all other proprioceptive observations, and the concatenated vector is then fed into an MLP with hidden layer sizes of [512, 256, 128] and ELU activation function. The base velocity estimator is a neural network that is used to achieve end-to-end mapping from sensor data to state quantities for speed estimation. The network for estimation, which is a MLP with hidden layer sizes [256, 128] and ELU activations, is trained in a supervised manner using ground-truth base velocity from simulator as labels.
@@ III-B Semantic-Aware Adaptive Foothold Planning
Building upon this kinematic prior, our proposed method performs a localized optimal search around the nominal Raibert position $\bm{P}_{\text{raibert}}^{i}$ leveraging the semantic grid map $\bm{H}_{t}^{\text{sem}}$. We construct a localized $M\times M$ search grid $\mathcal{G}^{i}$ with resolution $\Delta_{grid}$ centered around $\bm{P}_{raibert}^{i}$. For each candidate point $\bm{p}_{k}\in\mathcal{G}^{i}$, we transform it to the world frame and check for virtual collisions against known obstacles from the semantic module. To determine the safest and most balanced foothold, we design a cost function $J(\bm{p}_{k})$ composed of two terms: deviation cost and collision penalty.
[…]
where $\bm{p}_{k}-\bm{P}_{\text{raibert}}^{i}||_{2}$ is the deviation cost, which encourages the robot to select a foothold as close to the theoretically optimal Raibert point as possible to maintain stability. The collision penalty, $\mathbbm{1}_{\text{col}}(\bm{p}_{k})$, is an indicator function that evaluates to $1$ if the candidate point $\bm{p}_{k}$ triggers an Axis-Aligned Bounding Box (AABB) collision with any semantic obstacle, whose boundaries are uniformly dilated by the foot radius $r_{\text{foot}}$ to ensure a strict safety margin, and 0 otherwise. We assign a highly punitive weight $w_{\text{col}}\gg w_{\text{dev}}$ to ensure that unsafe locations are strictly avoided.
The final optimized target foothold $\bm{P}_{\text{target}}^{i}$ is derived by minimizing the cost function in all candidates:
@@ III-C Reward Functions
where $r_{\text{task}}\geq 0$ encourages the completion of primary objectives, $r_{\text{aux}}\leq 0$ penalizes undesirable behaviors, and $c_{\text{aux}}$ is a positive scaling factor. This form largely resolves the trade-off issue between different rewards. All reward terms are listed in Supplymentary Material A.
@@ III-C1 Primary Objective Rewards ($r_{\text{primary}}$)
In highly cluttered environments, tracking the user’s velocity command is insufficient for survival; precise foot placement is equally paramount. Therefore, our primary reward integrates both kinematic velocity tracking and semantic-aware foothold tracking:
@@ III-C2 State and Dynamic Constraint Penalties ($r_{\text{penalty}}$)
During the swing phase, the normalized swing progress $\phi_{\text{swing}}^{i}\in[0,1]$ for leg $i$ is calculated based on the stance duty factor. To synthesize an asymmetric safety-first swing profile characterized by rapid liftoff and a prolonged airborne plateau, we formulate the time-varying reference height $z_{\text{ref}}^{i}$ using a square-root sine function:
@@ III-D Learning Semantic-Aware Locomotion via Two-Stage RL
Learning semantic-aware locomotion from scratch in a highly cluttered environment poses a significant exploration challenge. Direct training with rigid physical obstacles frequently causes premature episode terminations (e.g., the robot tripping and falling), severely restricting the agent’s ability to collect meaningful reward signals and leading to overly conservative policies. Consequently, the policy often converges to suboptimal local minima, failing to master either stable locomotion or reliable obstacle avoidance simultaneously. To address this, we design a two-stage RL approach.
@@ III-D1 Virtual Obstacle Stage
In the first stage, we create a virtual training environment where obstacles are visually and semantically represented in $\bm{H}_{t}^{\text{elev}}$ and $\bm{H}_{t}^{\text{sem}}$, but physical collisions are disabled, as shown in Fig. 3. The robot’s limbs can pass through obstacles without triggering physical reaction forces or ending the episode. This means the robot will neither trip over obstacles nor push them away. During this stage, the policy relies solely on the Semantic-Aware Raibert Heuristic Reward ($r_{\text{sem}}$) to adjust its foot placement. Working together with the curriculum, this safe stage allows the policy to gradually progress from basic stable locomotion to semantic-aware foot placement. This helps the robot build a robust mapping from visual inputs to spatial leg-swing coordination, without the severe penalties of physical collisions.
Figure 3: Virtual semantic elevation map. Traversability are color-coded (red and yellow, respectively), with obstacle elevations represented by the vertical positions of the yellow spheres.
@@ III-D2 Rigid Obstacle Stage
After the policy converges in the soft dynamic stage, we transfer it to the second stage, where full physical interactions are enabled. Obstacles are created as rigid bodies, and their density and variety are increased compared to the first stage, further improving the robot’s obstacle avoidance ability. In this stage, the policy adjusts the previously learned motion patterns to handle complex physical interactions, such as friction changes and minor accidental collisions. This results in robust and collision-free locomotion that is suitable for real-world deployment.
@@ III-E2 Obstacle Density Curriculum
Figure 4: Example of obstacle curricula.
@@ III-F Training Setup in Simulator and Real-world Deployment
Training: We conduct two-stage training environments in the IsaacLab framework[21] and use Proximal Policy Optimization[27] to train policies. Each stage is trained with 4096 parallel environments of Unitree Go2 robot; the domain randomization ranges and the PPO hyperparameters. We train on a single NVIDIA RTX 5090 GPU, with the soft constraint stage taking approximately 1.5 hours and the hard constraint stage 3.5 hours.
Real-World Deployment: We deploy SemLoco on a Unitree Go2 equipped with a head-mounted spatial memory module Odin1 and an onboard computer. To preserve the robot’s nominal dynamics, all computations—including semantic segmentation, map fusion, and policy inference—are executed locally. The control policy runs at 50 Hz, while the asynchronous perception front-end updates at 8–10 Hz. Specifically, a semantic map, which is converted into a traversability cost map via class-specific weights and synchronized with a depth-derived elevation map. This dual-map representation ensures low-latency, foot-level perception. During testing, the robot navigates straight trajectories through indoor clutter (e.g., cables, scattered debris). For safety, forward speed is capped up to 0.7 m/s with automatic stopping under low map confidence, and all states are logged for post-run analysis.
@@ IV Experiments
Figure 5: Performance contrast among different policies in real world.
@@ IV-A Experimental Setup
  • Ours w/o Semantic Map: Our policy without semantic map by setting the semantic map to zero.
@@
Experiments are conducted in Isaac Sim [22]. For the simulation evaluation, we benchmark each policy on a $10\text{m}\times 2\text{m}$ straight-line track populated with varying densities of obstacle. In contrast to the obstacle placement strategy used during training, obstacles in the physical experiments may be stacked on top of each other, as shown in Supp.. The robot is commanded to traverse the track with a target forward velocity of $0.7\text{ m/s}$.
@@
  • Step Collision Rate ($C$, %): The percentage of footsteps that collide with obstacles relative to the total number of steps taken during traversal.

### v2

# Watch Your Step: Learning Semantically-Guided Locomotion in Cluttered Environments

@@ Abstract
Although legged robots demonstrate impressive mobility on rough terrain, using them safely in cluttered environments remains a challenge. A key issue is their inability to avoid stepping on low-lying objects, such as high-cost small devices or cables on flat ground. This limitation arises from a disconnection between high-level semantic understanding and low-level control, combined with errors in elevation maps during real-world operation. To address this, we introduce SemLoco, a Reinforcement Learning (RL) framework designed to avoid obstacles precisely in densely cluttered environments. SemLoco uses a two-stage RL approach that combines both soft and hard constraints. It performs pixel-wise foothold safety inference, which enables more accurate foot placement. Additionally, SemLoco integrates semantic map, allowing it to assign traversability costs instead of relying only on geometric data. SemLoco greatly reduces collisions and improves safety around sensitive objects, enabling reliable navigation in situations where traditional controllers would likely cause damage. Experimental results further show that SemLoco can be effectively applied to more complex, unstructured real-world environments. A demo video can be viewed at https://youtu.be/FSq-RSmIxOM.
Figure 1: SemLoco Overview for semantic-aware locomotion in cluttered environments. This figure demonstrates a quadruped robot performing obstacle avoidance in a real-world scene with small sensitive objects. Compared to the traditional pure elevation map, SemLoco integrates semantic map to low-level control, enabling the controller to perform pixel-wise foothold safety inference. This allows for precise gait planning in both simulation and real-world experiments, preventing damage to valuable items.
@@ I Introduction
  • SemLoco Framework: We propose the SemLoco framework, which solves the exploration trap in the navigation of dense obstacles. This decouples spatial exploration from strict dynamics, enabling robust perception-motor mappings without early-stage kinematic collapse.
  • Semantic-Geometry Decoupling: We introduce a control paradigm that explicitly disentangles environmental semantics from geometric elevation. This resolves the inherent vulnerabilities of traditional elevation-based proxies, making the system highly robust against the depth sensor noise and specular reflections typical of small, fragile objects.
  • Sim-to-Real Deployment and Validation: We rigorously validate the proposed method across both simulations and real-world scenarios featuring fragile equipment. Experiments demonstrate that SemLoco drastically reduces the step collision rate, proving its robust autonomy and adaptability in cluttered environment.
@@ II Related Works
Semantic Perception in Robot Navigation: Beyond purely geometry-based navigation, recent works leverage semantic information to enhance planning. Several modular approaches employ semantic segmentation models to obtain object masks and assign heuristic costs to construct dense semantic maps [4, 18] or semantic graphs [14]. [13] learns a semantic traversability estimator through an automated annotation pipeline on egocentric videos, simplifying the derivation of traversal cost maps. Furthermore, semantic-aware planners are built upon these representations to incorporate object-level risk into navigation decisions [3, 1]. More recently, [23] proposes an end-to-end path planning framework that encodes semantic information in the latent space, improving planning efficiency. [29] integrates semantic semantics into quadrupedal navigation, yet its impact is confined to the tuning of gait parameters, and velocity levels. [2] generates a passability cost map for local path planning by extracting rich image semantic features through the vision transformer DINOv2. These semantic-aware methods primarily operate at the high-level path planning stage and do not explicitly reason about foothold selection for legged robots. In cluttered and narrow environments, even if fragile or undesirable objects are recognized by the high-level planner, the absence of foothold-level constraints may still cause the robot to step on them, limiting the effectiveness of semantic awareness in contact-rich locomotion. To address the above issues, SemLoco   for the first time, combines semantic information with foothold planning for legged robots, using it to assign traversability costs to different types of obstacles, thereby establishing an effective connection between high-level semantic understanding and low-level motion control.
@@ III Methodology
Figure 2: Framework of SemLoco. Sub-modules have different styles based on their functions. Among them, the red trapezoid represents the neural network, the blue rectangle represents unprocessed raw data, and the green rectangle represents processed ready-to-use data. (a) Training in the simulator: In stage 1, we use virtual obstacles (Highly yellow spheres). Although the robot walks on flat ground, it receives a virtual perception map containing height and semantic information to simulate scenarios with real obstacles. In stage 2, we use rigid obstacles. The robot receives corresponding real perception information, and fine-tunes the policy to improve task performance. (b) Deployment in the real world: Exteroception information is obtained by Odin1 and fed into the semantic graph algorithm.
@@ III-A1 Observation Space
where $\bm{C}_{t}^{\text{cmd}}\in\mathbb{R}^{3}$ are the desired velocity commands, including the base linear velocities $\bm{v}_{x}^{\text{cmd}}\in\mathbb{R}^{1}$ in the x-axes, $\bm{v}_{y}^{\text{cmd}}\in\mathbb{R}^{1}$ in the y-axes and the base angular velocity $\bm{\omega}_{z}^{\text{cmd}}\in\mathbb{R}^{1}$ in the yaw axis.
@@
The $\bm{O}_{t}^{\text{extero}}\in\mathbb{R}^{1440}$ are defined as information about exteroception. Environmental information includes the elevation map $\bm{H}_{t}^{\text{elev}}\in\mathbb{R}^{720}$ and the semantic map $\bm{H}_{t}^{\text{sem}}\in\mathbb{R}^{720}$, both of which are centered at the center of gravity of the robot. Both maps are defined over a region that spans 1.5 m in the x-direction and 1.2 m in the y-direction. At 0.05m resolution, this results in a 30 $\times$ 24 grid of sampled cells.
@@ III-A3 Network Architecture
The framework is illustrated as Fig. 2. Our policy consists of two Multilayer Perceptrons (MLPs) for the actor network and the critic network respectively, a base velocity estimator and a Convolutional Neural Network (CNN) encoder. The CNN encoder extracts height features from elevation map and semantic features from semantic map. The feature vector extracted by the CNN module is concatenated with all other proprioceptive observations, and the concatenated vector is then fed into an MLP with hidden layer sizes of [512, 256, 128] and ELU activation function. The base velocity estimator, which is a MLP with hidden layer sizes [256, 128] and ELU activations, is trained in a supervised manner.
@@ III-B Semantic-Aware Adaptive Foothold Planning
Building upon this kinematic prior, our proposed method performs a localized optimal search around the nominal Raibert position $\bm{P}_{\text{raibert}}^{i}$ leveraging the semantic grid map $\bm{H}_{t}^{\text{sem}}$. We construct a localized $M\times M$ search grid $\mathcal{G}^{i}$ with resolution $\Delta_{\text{grid}}$ centered around $\bm{P}_{\text{raibert}}^{i}$. For each candidate point $\bm{p}_{k}\in\mathcal{G}^{i}$, we transform it to the world frame and check for virtual collisions against known obstacles from the semantic module. Crucially, rather than treating all spatial deviations equally, quadrupeds exhibit strong kinematic preferences during obstacle avoidance: shortening the stride (stepping backward relative to the nominal point) is biomechanically safer than over-stretching forward, while lateral deviations severely compromise the support polygon and should be strictly avoided. To capture this, we design a kinematically informed cost function $J(\bm{p}_{k})$ composed of an asymmetric directional deviation cost and a collision penalty:
$J(\bm{p}_{k})=C_{\text{dir}}(\bm{p}_{k})+w_{\text{col}}\cdot\mathbbm{1}_{\text{col}}(\bm{p}_{k}),$ (7)
where the directional deviation cost $C_{\text{dir}}(\bm{p}_{k})$ is defined in the robot’s local heading frame with longitudinal offset $\Delta x$ and lateral offset $\Delta y$:
$C_{\text{dir}}(\bm{p}_{k})=w_{x}^{+}\max(\Delta x,0)+w_{x}^{-}\max(-\Delta x,0)+w_{y}|\Delta y|.$ (8)
To enforce the aforementioned kinematic priors, we strictly set the penalty weights as $w_{y}\gg w_{x}^{+}>w_{x}^{-}$. The collision penalty, $\mathbbm{1}_{\text{col}}(\bm{p}_{k})$, is an indicator function that evaluates to $1$ if the candidate point $\bm{p}_{k}$ triggers an Axis-Aligned Bounding Box (AABB) collision with any semantic obstacle, whose boundaries are uniformly dilated by the foot radius $r_{\text{foot}}$ to ensure a strict safety margin, and 0 otherwise. We assign a highly punitive weight $w_{\text{col}}\gg w_{y}$ to ensure that unsafe locations are fundamentally vetoed.The final optimized target foothold $\bm{P}_{\text{target}}^{i}$ is derived by minimizing the cost function across all candidates:
@@
In this task, the semantic information is defined as ”fragility”(from 0 to 1). To obtain semantic information, we construct a semantic traversability map by projecting semantic predictions from RGB images into a global grid map. At time $t$, an open-vocabulary semantic segmentation model [28] predicts a class probability distribution $P(S_{t}=c|u,v)$ for each pixel $(u,v)$ in the RGB image. Each pixel is associated with a 3D point from the aligned point cloud and projected to a grid cell $g$ in the elevation map [6]. For each cell, we maintain a probability distribution over semantic classes $P_{t}(c|g)$, which is updated using Bayesian fusion:
$P_{t}(c|g)\propto P(S_{t}=c|p_{i})P_{t-1}(c|g),$ (10)
where $p_{i}$ is the observed point projected to cell $g$. Each class $c$ is assigned a predefined traversal cost $\phi(c)$, and the final traversability cost of the cell is computed as the expected cost:
$M(g)=\sum_{c\in\mathcal{C}}P_{t}(c|g)\phi(c).$ (11)
@@ III-C Reward Functions
where $r_{\text{task}}\geq 0$ encourages the completion of primary objectives, $r_{\text{aux}}\leq 0$ penalizes undesirable behaviors, and $c_{\text{aux}}$ is a positive scaling factor. This form largely resolves the trade-off issue between different rewards. The terms in (12) are further elaborated below.
@@ III-C1 Primary Objective Rewards ($r_{\text{primary}}$)
In highly cluttered environments, tracking the velocity command is insufficient for survival, and precise foot placement is equally important. Thus, our primary reward integrates both kinematic velocity tracking and semantic-aware foothold tracking:
@@ III-C2 State and Dynamic Constraint Penalties ($r_{\text{penalty}}$)
During the swing phase, the normalized swing progress $\phi_{\text{swing}}^{i}\in[0,1]$ for leg $i$ is calculated based on the stance duty factor. To synthesize an asymmetric safety-first swing profile characterized by rapid liftoff and a prolonged airborne plateau (Fig. 3), we formulate the time-varying reference height $z_{\text{ref}}^{i}$ using a square-root sine function:
@@ III-D Learning Semantic-Aware Locomotion via Two-Stage RL
Learning semantic-aware locomotion from scratch in a highly cluttered environment poses a significant exploration challenge. Direct training with rigid physical obstacles frequently causes a large number of penalties, severely restricting the agent’s ability to collect meaningful reward signals and leading to overly conservative policies. Consequently, the policy often converges to suboptimal local minima, failing to master either stable locomotion or reliable obstacle avoidance. To address this, we design a two-stage RL approach.
@@ III-D1 Virtual Obstacle Stage
In the first stage, we create a virtual training environment where obstacles are visually and semantically represented in $\bm{H}_{t}^{\text{elev}}$ and $\bm{H}_{t}^{\text{sem}}$, but physical collisions are disabled, as shown in Fig. 2(a). The robot’s limbs can pass through obstacles without triggering physical reaction forces or ending the episode. This means the robot will neither trip over obstacles nor push them away. During this stage, the policy relies solely on the Semantic-Aware Raibert Heuristic Reward ($r_{\text{sem}}$) to adjust its foot placement. Working together with the curriculum, this safe stage allows the policy to gradually progress from basic stable locomotion to semantic-aware foot placement. This helps the robot build a robust mapping from visual inputs to spatial leg-swing coordination, without the severe penalties of physical collisions.
Figure 3: Footswing tracking for each leg.
@@ III-D2 Rigid Obstacle Stage
After the policy converges in the soft dynamic stage, we transfer it to the second stage, where full physical interactions are enabled. Obstacles are created as rigid bodies, and their density and variety are increased compared to the first stage, further improving the robot’s obstacle avoidance ability. In this stage, the policy adjusts the previously learned motion patterns to handle complex physical interactions, such as friction.
@@ III-E2 Obstacle Density Curriculum
Figure 4: Examples of obstacle curricula.
@@ III-F Training Setup in Simulator and Real-world Deployment
Training: We conduct two-stage training environments in the IsaacLab framework [21] and use Proximal Policy Optimization [26] to train policies. Each stage is trained with 4096 parallel environments of Unitree Go2 robot. We train on a single NVIDIA RTX 5090 GPU, with the soft constraint stage taking 1.5 hours and the hard 3.5 hours.
Real-World Deployment: We deploy SemLoco on a Unitree Go2 equipped with a head-mounted RGB-ToF module Odin1 and an AGX Orin. To preserve the robot’s nominal dynamics, all computations—including semantic segmentation, map fusion, and policy inference—are executed locally. The control policy runs at 50 Hz, while the asynchronous perception front-end updates at 10 Hz.
@@ IV Experiments
TABLE I: Ablation experiment under different obstacle densities.
(m) (%) (%) (m) (%) (%) (m) (%) (%) (m) (%) (%)
Ours 9.67 97.00 2.37 9.51 92.60 8.78 9.29 87.40 13.45 8.94 82.20 24.33
Ours w/o Virtual Obstacle Stage 6.19 51.60 13.18 5.88 45.80 20.88 5.42 39.40 29.67 5.06 30.80 36.15
@@ IV-A Experimental Setup
  • Ours w/o Semantic Map: Search using high-level information instead of semantic information.
@@
Experiments are conducted in Isaac Sim. For the simulation evaluation, we benchmark each policy on a $10\text{m}\times 2\text{m}$ straight-line track populated with varying densities of obstacle. The robot is commanded to traverse the track with a target forward velocity of $0.4\text{ m/s}$.
@@
  • Step Collision Rate ($C$, %): The percentage of footsteps that collide with obstacles relative to the total number of steps taken during the traversal.